### v4

# RL-MILP Solver: A Reinforcement Learning Approach for Solving Mixed-Integer Linear Programs with Graph Neural Networks

Abstract
Mixed-integer linear programming (MILP) is a widely used optimization technique across various fields. Existing end-to-end learning methods for MILP generate values for a subset of decision variables and delegate the remaining problem to traditional MILP solvers. However, this approach often fails to guarantee solution feasibility (i.e., satisfying all constraints) due to inaccurate predictions and primarily focuses on binary decision variables. Satisfying all constraints is a prerequisite for obtaining the optimal solution, and the feasibility issue becomes even more critical with non-binary integer (integer, for short) variables. Thus, addressing the feasibility of MILP involving integer variables is crucial. To address these challenges, we propose a novel reinforcement learning (RL)-based solver that not only finds the first feasible solution but also incrementally discovers better feasible solutions without delegating the remainder to off-the-shelf solvers. Our experimental results demonstrate that the proposed method achieves (near-)optimal solutions.
Reinforcement Learning, MILP, Graph Neural Networks, Combinatorial Optimization, Machine Learning, ICML
11footnotetext: School of Computing, Korea Advanced Institute of Science and Technology, Daejeon, Republic of Korea.
Corresponding author: Min-Soo Kim $<$minsoo.k@kaist.ac.kr$>$.
1 Introduction
The traveling salesman problem and the knapsack problem are classic examples of combinatorial optimization (CO) problems, extensively studied in operations research and computer science (Gasse et al., 2022). CO involves mathematical optimization that aims to minimize or maximize a specific objective function (Mazyavkina et al., 2021). If the objective function and constraints of CO are linear, it is called linear programming (LP). Furthermore, if some variables in LP must take integer values, it becomes Mixed-integer linear programming (MILP) (Bengio et al., 2021). MILP is widely applied to real-world scenarios such as logistics (Kweon et al., 2024) and path planning (Zuo et al., 2020).
ML methods that directly learn and generate solutions for CO problems are categorized as an end-to-end learning approach (Bengio et al., 2021). ML-based primal heuristics are considered as end-to-end learning (Han et al., 2023), which aim to quickly find solutions to MILP by reducing the search space (Achterberg et al., 2012). Unlike traditional primal heuristics, which rely on expert knowledge and hand-crafted designs (Bengio et al., 2021), the ML-based one leverages similar patterns shared by MILP instances generated from a specific distribution (Gasse et al., 2022). Figure 1(a) illustrates how ML-based primal heuristics solve MILP. A trained ML model generates a partial solution for a subset of integer variables, which is then passed to a traditional solver (e.g., Gurobi and SCIP) to optimize the remaining sub-problem, referred to as partial end-to-end learning.
(a) Partial end-to-end learning.
(b) Full end-to-end learning (Ours).
Figure 1: Solving MILP with end-to-end learning methods.
Existing ML-based primal heuristics have demonstrated their ability to quickly find good solutions (Ding et al., 2020; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024). However, inaccurate ML predictions hinder solution feasibility, which emphasizes the importance of feasibility (Han et al., 2023). The infeasibility obstructs reaching the optimal solution, posing a significant obstacle to solving MILP. Moreover, the existing studies mostly focus on the prediction for binary decision variables. However, many real-world problems involve non-binary integer (integer, for short) variables such as logistics (Kweon et al., 2024), maritime transportation (Papageorgiou et al., 2014), and energy systems (Ren & Gao, 2010). This highlights the need for techniques that can handle integer variables.
Infeasibility caused by incorrect ML predictions can be more pronounced for integer variables due to their broader range than binary variables. Imagine a naïve ML approach that predicts the value of an integer variable $x_{i}$ whose range is $[0,1000]$. The maximum possible prediction error for $x_{i}$ would be larger than binary variables because the range of $x_{i}$ is broader than binary variables whose range is $[0,1]$. Moreover, rounding predicted values to integers may cause additional accuracy loss (Cont & Heidari, 2014). A method for handling integer variables in partial end-to-end learning has been proposed, representing values in binary format. (Nair et al., 2020). In binary format, $x_{i}$ has the sequence length of $\left\lceil\log_{2}(1000)\right\rceil=10$, requiring multiple dimensions for a single variable. Even in some cases, the upper/lower bounds of variables are infinite, as seen in the default setting in SOTA solver, Gurobi Optimization. Thus, it calls for a more sophisticated and scalable approach (Cantürk et al., 2024).
To address this, we propose a RL-based method for solving MILP, called RL-MILP solver. Figure 1(b) illustrates how our solver derives the final solution. Unlike the partial end-to-end learning methods that focus on a subset of binary variables, the RL-MILP solver generates complete feasible solutions for MILP involving integer variables without delegating the sub-problem to traditional solvers. Moreover, as a RL-based method, the RL-MILP solver does not require the labeled data in contrast to those methods. As an initial study on full end-to-end RL-based MILP solvers, our study opens the door to a unified LP solver for multiple CO tasks, which is desirable in real-world scenarios (Liu et al., 2024).
We design a RL system tailored to MILP, which enables the agent to learn the variable-constraint relationships. Our method decides whether to change the value of integer variables rather than directly predicting their exact values. The solution search process is divided into two phases: before and after finding the first feasible solution. We design a reward function to align with the objectives of each phase. The RL agent is trained based on the degree of constraint violations and the solution quality. We adopt a Transformer-based GNN to capture relationships among distant variables. The agent improves the solution quality guided by the proposed search strategy. Experimental results demonstrate that the RL-MILP solver achieves optimal solutions on a small dataset and finds near-optimal solutions under a 1% gap for larger datasets. Our main contributions are as follows:
  • We propose a novel RL-MILP solver that incrementally improve feasible solutions for MILP involving integer variables without reliance on off-the-shelf solvers.
  • We design a RL system that enables the agent to learn problem-solving capabilities by capturing relationships between decision variables and constraints.
  • We propose learning and search strategies for solving MILP problems grounded in a theoretical foundation.
  • We propose a Transformer-based GNN as the agent’s architecture, designed to effectively capture the complex relationships among decision variables.
2 Preliminaries
2.1 Mixed-Integer Linear Programming
2.1.1 A Standard Form
Mixed-integer linear programming (MILP) is an optimization problem that minimizes or maximizes a linear objective function while satisfying linear constraints and integrality requirements for some decision variables (Bertsimas & Tsitsiklis, 1997). A standard form of MILP is as follows:
minimize $\mathbf{c}^{T}\mathbf{x}$ (1)
subject to $\mathbf{A}\mathbf{x}\leq\mathbf{b}$ (2)
$x_{i}\in\mathbb{Z},\forall i\in I$ (3)
$l_{i}\leq x_{i}\leq u_{i},\forall i$ (4)
where $\mathbf{x}\in\mathbb{R}^{n}$ is a column vector of $n$ decision variables, $\mathbf{c}\in\mathbb{R}^{n}$ is a column vector of the objective coefficients, $\mathbf{A}\in\mathbb{R}^{m\times n}$ is the constraint coefficient matrix, $\mathbf{b}\in\mathbb{R}^{m}$ is a column vector of the right-hand side of the constraints, $I$ is the index set of integer decision variables, $l_{i}$/$u_{i}$ denotes the lower/upper bounds for each decision variable $x_{i}$. The goal of MILP is to find the optimal solution, and for a minimization problem, this corresponds to a feasible solution $\mathbf{x}$ that minimizes $obj=\mathbf{c}^{T}\mathbf{x}$ (Eq. 1). A feasible solution is a solution $\mathbf{x}$ that satisfies all constraints (Eqs. 2-4).
All MILP problems can be transformed into the standard form (Eqs. 1-4) (Bertsimas & Tsitsiklis, 1997). Let $\mathbf{a}_{i}^{T}$ denote a row vector of a single constraint, $\mathbf{A}=\left(\mathbf{a}_{1}^{T},\ldots,\mathbf{a}_{m}^{T}\right)$, and $\mathbf{b}=\left(b_{1},\ldots,b_{m}\right)$. An equality constraint $\mathbf{a}_{i}^{T}\mathbf{x}=b_{i}$ is equivalent to two inequality constraints ($\mathbf{a}_{i}^{T}\mathbf{x}\geq b_{i}$ and $\mathbf{a}_{i}^{T}\mathbf{x}\leq b_{i}$). Moreover, $\mathbf{a}_{i}^{T}\mathbf{x}\geq b_{i}$ is equivalent to $-\mathbf{a}_{i}^{T}\mathbf{x}\leq-b_{i}$. Similarly, maximizing $\mathbf{c}^{T}\mathbf{x}$ is equivalent to minimizing $-\mathbf{c}^{T}\mathbf{x}$. Regardless of the direction of the objective function or constraints, a problem can be reformulated in the standard form. Thus, we only address standard-form problems (i.e., minimization) in the following sections.
2.1.2 Key Properties of MILP
MILP is known to be NP-hard, with integrality requirements (Eq. 3) (Nair et al., 2020). As the number of integer variables increases, its computation cost grows exponentially (Floudas, 1995). LP-relaxation refers to the technique of removing the integrality requirement, which transforms an original MILP problem into an LP problem solvable in polynomial time (Bertsimas & Tsitsiklis, 1997). LP-relaxation is commonly used to obtain an initial solution and a lower bound for the original problem (Witzig et al., 2021) as such in traditional algorithms (e.g., Branch-and-Bound (Land & Doig, 2010) and Feasible Pump (Fischetti et al., 2005)) Letting $obj_{r}$ be the optimal value of the LP-relaxed problem and $obj_{b}$ be the objective value obtained by the current best feasible solution of the given original MILP problem, Proposition 1 is derived as follows:
Proposition 1.
If a better feasible solution $\mathbf{x}^{\prime}$ exists for a given MILP problem, $\mathbf{x}^{\prime}$ must yield $obj^{\prime}$ that lies between $obj_{r}$ and $obj_{b}$.
The proof of Proposition 1 is provided in Appendix A. Drawing on Proposition 1, we design our RL system to guide exploration toward the optimal solution.
2.2 Graph for MILP and Graph Neural Networks
2.2.1 Bipartite Graph Representation of MILP
Studies on ML-based primal heuristics represent MILP instances as bipartite graphs (Ding et al., 2020; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024). In a bipartite graph representation of MILP, one set of nodes is for constraints, and the other is for decision variables. An edge connects a variable node to a constraint node only if the variable appears in the corresponding constraint. For example, in Figure 2(a), the variable $x_{3}$ appears in the constraint $\mathbf{a_{2}}$. Therefore, in the MILP bipartite graph, the variable node for $x_{3}$ is connected to the constraint node for $\mathbf{a_{2}}$.
2.2.2 Message Passing Neural Network
The message passing neural network (MPNN) (Gilmer et al., 2017) is a general framework for message passing-based GNNs. Widely used GNNs such as GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), and GIN (Xu et al., 2019) are the architectures based on MPNN. Given a graph $G$, the new representation $h_{v}^{(k+1)}$ is obtained as follows:
$m_{vu}^{(k+1)}=msg(h_{v}^{(k)},h_{u}^{(k)},e_{vu}),\forall e_{vu}\in\mathcal{E}$
$g_{v}^{(k+1)}=agg(\{m_{vu}^{(k+1)}|u\in\mathcal{N}(v)\}),\forall v\in\mathcal{V}$ (5)
$h_{v}^{(k+1)}=update(h_{v}^{(k)},g_{v}^{(k+1)}),\forall v\in\mathcal{V}$
where $h_{v}^{(0)}$ is the initial feature vector of a target node $v$, and $msg(\cdot)$, $agg(\cdot)$, and $update(\cdot)$ are the message-passing, aggregation, and update functions, respectively. $msg(\cdot)$ generates a message $m_{vu}^{(k+1)}$ using the representation of the target node $h_{v}^{(k)}$, the neighbor node $h_{u}^{(k)}$, and the edge feature $e_{vu}$. $agg(\cdot)$ aggregates the messages $m_{vu}^{(k+1)}$ generated by $msg(\cdot)$ for each target node. $update(\cdot)$ updates the target node $v$ to a new representation $h_{v}^{(k+1)}$ by combining the aggregated information $g_{v}^{(k+1)}$ with the previous embedding $h_{v}^{(k)}$. The embedding $h_{v}^{(k+1)}$ is used for prediction tasks or message passing in the next layer.
Since edges exist only between nodes from different sets (i.e., variable-constraint) in a MILP bipartite graph, obtaining a new representation for a decision variable requires two rounds of message passing. As shown in Figure 2(b), a new representation for the decision variables is obtained by performing one constraint-side convolution and one variable-side convolution after applying initial embeddings.
(a) Example of a bipartite graph representation for a MILP instance with three variables and two constraints.
(b) GCN for bipartite graph representation of MILP. V: Variable nodes, C: Constraint nodes, E: Edge features and adjacency matrix.
Figure 2: Bipartite graph for MILP and its GCN layers.
2.2.3 Transformer for Graphs
The Transformer (Vaswani, 2017), originally designed as a sequence-to-sequence model for machine translation, has achieved significant success in various domains such as NLP and computer vision. Recently, there have been numerous efforts to adapt Transformers for the graph domain (Lin et al., 2022; Min et al., 2022). The attention mechanism of Transformers enables each node to attend to every node, which allows the model to effectively learn relationships between distant nodes (Wu et al., 2021).
MPNN-based GNNs (MPNN, for short) receive messages from the neighbor node, which is suitable for learning local structural information. However, they struggle to capture relationships between distant nodes (Zhang et al., 2020; Wu et al., 2021). To propagate messages from nodes that are $K$ hops away, a MPNN requires $K$ layers. However, deeper layers can lead to oversmoothing, where nodes in a graph have similar and indistinguishable representations (Li et al., 2018; Wu et al., 2021; Min et al., 2022).
A MPNN for MILP may require deep layers to capture the relationships among variables that influence each other across multiple constraints. For instance, as illustrated in Figure 2(a), the variables $x_{2}$ and $x_{3}$ are 4 hops apart ($x_{2}$ $\mathbf{a_{1}}$ $x_{1}$ $\mathbf{a_{2}}$ $x_{3}$). Although $x_{2}$ and $x_{3}$ do not appear in the same constraint, they are connected via $x_{1}$, which appears in both $\mathbf{a_{1}}$ and $\mathbf{a_{2}}$. A change in the value of $x_{2}$ can affect $x_{1}$, which may subsequently affect $x_{3}$. Thus, capturing the relationship between these variables is essential. Propagating messages from $x_{2}$ to $x_{3}$ requires four GCN layers (see Figure 2). However, even a shallow MPNN with 2-4 layers can suffer from the oversmoothing (Wu et al., 2023). To address this, we utilize a Transformer-based GNN, which can effectively learn relationships among variables, regardless of their distance.
Figure 3: The overview of our approach. The symbols $\uparrow$, $\downarrow$, and -- denote an increase, a decrease, and no change in the values of decision variables, respectively. The white point ($\mathbf{x}_{0}$), red point ($\mathbf{x}_{1}$), and green points ($\mathbf{x}_{2}$, $\mathbf{x}_{3}$) denote the initial solution, an infeasible solution, and feasible solutions, respectively. The darker the area, the better the solutions.
3 Methodology
This section introduces the RL-MILP solver in detail. Figure 3 illustrates how our solver incrementally improves feasible solutions. The process begins with converting a given MILP instance into a bipartite graph. Next, the agent is trained on the graph to select high-reward actions. The selected actions derive a new solution $\mathbf{x}_{t+1}$. Finally, the current best solution $\mathbf{x}_{b}$ is updated if $\mathbf{x}_{t+1}$ is feasible and better than $\mathbf{x}_{b}$.
3.1 Reinforcement Learning for MILP
A RL system for MILP aims to train the agent to make decisions that maximize rewards while interacting with a given instance. Figure 4 depicts how the RL agent interacts with a MILP instance, where $\mathcal{S}_{t}$, $\mathcal{A}_{t}$, and $\mathcal{R}_{t,total}$ denote the observation, the set of selected actions, and the total reward at timestep $t$, respectively. The instance $M$ acts as the environment for the agent. Using $\mathcal{A}_{t}=(a_{t,1},\dots,a_{t,n})$, the agent updates the solution $\mathbf{x}_{t+1}$ for $n$ variables. This update affects the left-hand side of the constraints $\mathbf{lhs}_{t+1}$, the feasible state vector $\mathbf{f}_{t+1}$, and the objective value $obj_{t+1}$. At the next timestep, the agent faces a new observation $\mathcal{S}_{t+1}$ changed by its previous actions and selects a new action set $\mathcal{A}_{t+1}$ to maximize rewards. By comparing the actual reward $\mathcal{R}_{t+1,total}$ obtained from $\mathcal{A}_{t+1}$ with the estimated reward, the agent refines its policy $\pi$.
Figure 4: Diagram of reinforcement learning for MILP.
3.1.1 Observation
We define the observation $\mathcal{S}_{t}=(\mathbf{x}_{t},\mathbf{f}_{t},obj_{t})$. The solution $\mathbf{x}_{t}$ is derived by updating the values of variables based on the agent’s actions $\mathcal{A}_{t-1}$. For example, assuming $\mathcal{A}_{t-1}=(a_{t-1,1},\,a_{t-1,2},\,a_{t-1,3})=(+1,+0,+0)$, then $\mathbf{x}_{t-1}$ in Figure 4 is updated to $\mathbf{x}_{t}=(x_{t,1},\,x_{t,2},\,x_{t,3})=(3,9,-2)$. With the updated $\mathbf{x}_{t}$, the new $\mathbf{lhs}_{t}=\mathbf{A}\mathbf{x}_{t}$ and $obj_{t}=\mathbf{c}^{T}\mathbf{x}_{t}$ are calculated.
Each elements of $\mathbf{f}_{t}$ indicates whether $\mathbf{x}_{t}$ satisfies the corresponding constraint, defined as $\mathbf{f}_{t}=\mathbf{b}-\mathbf{lhs}_{t}$. Non-negative elements in $\mathbf{f}_{t}$ indicate satisfied constraints, while negative ones indicate violations. For example, in Figure 4, $\mathbf{x}_{t+1}$ yields $\mathbf{lhs}_{t+1}=(31,4)$ for constraints $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$. As $\mathbf{f}_{t+1}=\mathbf{b}-\mathbf{lhs}_{t+1}=(30,5)-(31,4)=(-1,1)$, $\mathbf{x}_{t+1}$ violates $\mathbf{a}_{1}$ but satisfies $\mathbf{a}_{2}$.
3.1.2 Action
At timestep $t$, the agent selects a set of actions $\mathcal{A}_{t}=(a_{t,1},\dots,a_{t,n})$ for $n$ variables based on $\mathcal{S}_{t}$. For each variable, the agent can take one of three types of actions: increase, decrease, or no change, as shown in Figure 4.
3.1.3 Reward
We design reward functions that enable the agent to select actions that maximize the total reward $\mathcal{R}_{t,\text{total}}$ in a given situation, as follows:
$\mathcal{R}_{t,\text{total}}=\mathcal{R}_{t,\text{opt}}+\mathcal{R}_{t,\text{explore}}$ (6)
Finding a feasible solution is a prerequisite for discovering the optimal solution. Therefore, the agent’s primary objective is to find a feasible solution that satisfies all constraints. The feasibility reward $\mathcal{R}_{t,\text{F}}$ is calculated based on the constraint satisfaction and the upper/lower bound conditions of the decision variables, as follows:
$\mathcal{R}_{t,\text{F}}=\frac{1}{\sqrt{mn}}\mathcal{R}_{t,\text{const}}+\frac{1}{\sqrt{n}}\left(\mathcal{R}_{t,\text{ub}}+\mathcal{R}_{t,\text{lb}}\right)$ (7)
$\mathcal{R}_{t,\text{const}}=\sum_{j=1}^{m}\min(f_{t+1,j},0)-\min(f_{t,j},0)$ (8)
$\mathcal{R}_{t,\text{ub}}=\sum_{i=1}^{n}\min(u_{i}-x_{t+1,i},0)-\min(u_{i}-x_{t,i},0)$ (9)
$\mathcal{R}_{t,\text{lb}}=\sum_{i=1}^{n}\min(x_{t+1,i}-l_{i},0)-\min(x_{t,i}-l_{i},0)$ (10)
where $f_{t,j}$ is the element of $\mathbf{f}_{t}$ for the $j$-th constraint, and $x_{t,i}$ of $\mathbf{x}_{t}$ is the value of the $i$-th decision variable at timestep $t$. The constraint reward $\mathcal{R}_{t,\text{const}}$ is proportional to improvements (or deteriorations) for each infeasible constraint. For example, in Figure 4, $\mathcal{R}_{t,\text{const}}$ is [$[\{-1-(-3)\}+\{0-(-2)\}]=4$. The bound rewards $\mathcal{R}_{t,\text{ub}}$ and $\mathcal{R}_{t,\text{lb}}$ are proportional to changes for each variable that violates its bound conditions. As shown in Figure 4, $\mathcal{R}_{t,\text{lb}}$ is [{${}[\{0-0\}+\{0-0\}+\{-1-(-2)\}]=1$ because the lower bound is 0, and $x_{t+1,3}$ improves by 1.
The reward system for our solver operates in two phases: $phase1$ and $phase2$. $phase1$ continues until the first feasible solution is found, while $phase2$ begins thereafter. The optimization reward $\mathcal{R}_{t,\text{opt}}$ is calculated, as follows:
$\mathcal{R}_{t,\text{opt}}=\begin{cases}\mathcal{R}_{t,\text{p1}},&\text{if agent is in $phase1$},\\ \mathcal{R}_{t,\text{p2}},&\text{otherwise}.\end{cases}$ (11)
$\mathcal{R}_{t,\text{p1}}=\begin{cases}1-\delta_{t,r},&\text{if }\mathbf{x}_{t+1}\in\mathcal{F},\\ \mathcal{R}_{t,\text{F}},&\text{otherwise}.\end{cases}$ (12)
$\mathcal{R}_{t,\text{p2}}=\begin{cases}(1-\delta_{t,r})\cdot\alpha,&\text{if }\mathbf{x}_{t+1}\in\mathcal{F}\land obj_{t+1}<obj_{b},\\ 0,&\text{if }\mathbf{x}_{t+1}\in\mathcal{F}\land obj_{t+1}\geq obj_{b},\\ \mathcal{R}_{t,\text{F}}\cdot\delta_{t,b}\cdot\beta,&\text{if }\mathbf{x}_{t+1}\notin\mathcal{F}\land obj_{t+1}<obj_{b},\\ \mathcal{R}_{t,\text{F}}\cdot\delta_{t,b}&\text{otherwise}.\end{cases}$ (13)
$\delta_{t,r}=\frac{|obj_{t+1}-obj_{r}|}{|obj_{r}|},\quad\delta_{t,b}=\frac{|obj_{t+1}-obj_{b}|}{|obj_{r}-obj_{b}|}.$ (14)
where $\mathcal{F}$ is the feasible region, $\alpha=1/\beta$ and $\beta\in(0,1]$ are toward-optimal bias parameters, and $\delta_{t,r}$ and $\delta_{t,b}$ are scaled gaps from $obj_{t+1}$ to the lower bound $obj_{r}$ and the current best objective value $obj_{b}$, respectively.
In $phase1$, the agent aims to find the first feasible solution. Achieving a better (i.e., lower) objective value in $phase1$ sets the agent up for a better starting point in $phase2$. Thus, higher rewards are given as the solution nears $obj_{r}$ (Eq. 12). If the agent finds an infeasible solution $\mathbf{x}_{t+1}\notin\mathcal{F}$, it is corrected by the feasibility reward $\mathcal{R}_{t,\text{F}}$. In $phase2$, the goal is to find feasible solutions better than the current best solution $\mathbf{x}_{b}$. If the agent finds better $\mathbf{x}_{t+1}\in\mathcal{F}$, it earns higher rewards as $\mathbf{x}_{t+1}$ gets closer to $obj_{r}$ (Case 1 in Eq. 13). For $\mathbf{x}_{t+1}\in\mathcal{F}$ worse than $\mathbf{x}_{b}$, the agent receives no reward due to no improvement and is also not penalized to avoid invalidating the learning from in $phase1$ (Case 2 in Eq. 13). If $\mathbf{x}_{t+1}\notin\mathcal{F}$, the agent receives smaller $\mathcal{R}_{t,\text{F}}$ as it gets closer to $obj_{b}$ (Case 3, 4 in Eq. 13). According to Proposition 1, we regard $obj_{t+1}<obj_{b}$ as desirable; thus, the penalty is reduced by $\beta$ (Case 3 in Eq. 13). Adjusting $\beta$ further promotes exploration toward $obj_{r}$. Appendix B provides further visual explanations of the reward functions.
To promote exploration, $\mathcal{R}_{t,\text{explore}}$ is calculated as follows:
$\mathcal{R}_{t,\text{explore}}=\begin{cases}-10,&\text{if }\mathbf{x}_{t+1}=\mathbf{x}_{t},\\ 0,&\text{otherwise}.\end{cases}$ (15)
To prevent the agent from ceasing exploration, which is highly undesirable, the highest penalty of -10 is imposed.
3.1.4 Learning Algorithm
We adopt the Actor-Critic (AC) algorithm (Mnih et al., 2016), which has demonstrated effectiveness in solving CO problems (Bello et al., 2016; Hubbs et al., 2020; Zong et al., 2022). AC combines a policy-based actor and a value-based critic. The actor optimizes parameters $\theta$ to maximize the expected reward by learning a policy $\pi_{\theta}(\mathcal{A}\mid\mathcal{S})$ that maps the observation $\mathcal{S}$ to a probability distribution over actions $\mathcal{A}$. The critic evaluates $\mathcal{S}$ by learning a value function $V_{\theta}(\mathcal{S})$, which estimates expected return for a given $\mathcal{S}$.
At each timestep $t$, the agent observes $\mathcal{S}_{t}$ and selects actions $\mathcal{A}_{t}$ using the policy $\pi$. Upon executing $\mathcal{A}_{t}$, the environment reacts by providing a observed reward $\mathcal{R}_{t,\text{total}}$ and $\mathcal{S}_{t+1}$. The actor is trained to encourage actions with higher $\mathcal{R}_{t,\text{total}}$ than the estimated value $V_{\theta}(\mathcal{S}_{t})$ and suppress those with lower $\mathcal{R}_{t,\text{total}}$. The critic is trained to minimize the gap between $\mathcal{R}_{t}$ and $V_{\theta}(\mathcal{S}_{t})$ to provide accurate feedback to the actor.
To train the RL agent across diverse environments, we expose it to randomly generated MILP instances. To ensure sufficient training in $phase1$, we enforce the agent stays in $phase1$ for the predefined number of steps, even after finding the first feasible solution. Training proceeds until the agent reaches the total step limit, finds the optimal solution, or fails to improve solutions in $phase2$ within the grace period. Once any condition is met, the agent switches to a new instance. Appendix C presents the pseudo-code.
3.2 GNN Architecture of RL Agent
We adopt a hybrid approach combining a MPNN with a Transformer encoder (Rong et al., 2020; Wu et al., 2021) for the RL agent, as shown in Figure 5. The MPNN captures local information for each decision variable. Local information can include whether the constraints involving a variable are satisfied. To learn local information, we use the following features: $\mathbf{x}_{t}$ for variable nodes, $\mathbf{b}$ for constraint nodes, and $\mathbf{A}$ for edges. To ensure stable training, we scale $\mathbf{A}$ to $[-1,1]$ using equilibration scaling (Tomlin, 1975), which normalizes each constraint by its largest absolute coefficient. This scaling relies on the principle that multiplying each constraint by a positive scalar produces an equivalent LP problem to the original. For $\mathbf{x}_{t}$ and $\mathbf{b}$, which may have unbounded values, we use a Periodic Embedding (PE) (Gorishniy et al., 2022) module to embed scalars into vectors. PE has demonstrated effectiveness in deep learning models for processing numerical features, including tasks such as house price and income prediction (Pace & Barry, 1997; Kohavi et al., 1996). PE is formulated as follows:
$\operatorname{PE}(z)=\oplus(\sin(\tilde{z}),\cos(\tilde{z})),\;\tilde{z}=\left[2\pi w_{1}z,\ldots,2\pi w_{k}z\right].$
where ($\oplus$($\cdot$)⊕ ( ⋅ ) is concatenation, with scalar $z$ and trainable $w_{i}$.
[…]
The Actor layer generates $\mathcal{A}_{t}$ using the final embeddings for decision variables. The Critic layer approximates the expected return function using the embeddings for [PHA], $obj_{r}$, $obj_{b}$, and $\mathbf{f}_{t}$. These embeddings are used because the total reward depends on the solution quality and feasibility improvements or deteriorations (see Eq. 6). Since the reward system differs between $phase1$ and $phase2$, the Actor and Critic layers are separated by phase, while all other layers share parameters across the two phases.
3.3 Solution Search Strategy
Our search strategy is inspired by local search (LS), a widely used heuristic that explores the neighbors of the current solution for CO problems (Bertsimas & Tsitsiklis, 1997; Hillier & Lieberman, 2015). While LS may encounter local optima, empirical evidence suggests that it is effective at quickly finding quality solutions (Bertsimas & Tsitsiklis, 1997). Classic LS randomly tweaks one variable at a time, but moving farther in each step can be more effective (Shaw, 1998). In contrast to the classic LS, our search strategy imposes no restrictions on the number of variables that can change per step. Thus, the RL-MILP solver explores a larger neighborhood than LS conceptually. Moreover, the magnitude of the changes can be adjusted by expanding the number of the agent’s possible actions. For instance, each variable can have three possible actions (+1, 0, -1) or five (+2, +1, 0, -1, -2). Our approach employs a trained RL agent to navigate the solution space toward the optimal solution. The agent selects actions expected to yield higher rewards based on new observations. During $phase1$, the agent searches within a local region, whereas in $phase2$, it moves only to feasible solutions with better or equal objective values. Appendix D provides the pseudo-code.
Figure 5: The GNN architecture for solving MILP.
4 Experiments
We conduct experiments to answer these research questions:
  • RQ1: How accurately does RL-MILP solver approximate solutions in terms of solution quality?
  • RQ2: How does the toward-optimal bias parameter $\beta$ influence the solution quality?
  • RQ3: How effective is the proposed GNN architecture compared to alternative architectures in solving MILP?
4.1 Experimental Setup
4.1.1 Dataset Collection
We generated MILP instances based on the description in (Qi et al., 2021). Our study focuses on a special case of MILP where all variables are integers to emphasize handling integer variables. Datasets are named $D_{n\times m}$, where $n$ and $m$ are the numbers of variables and constraints, respectively. We use $D_{9\times 18}$, $D_{50\times 20}$, $D_{100\times 50}$, and $D_{300\times 100}$, where $D_{300\times 100}$ is approximately 185 times larger than $D_{9\times 18}$ that matches the largest dataset from (Qi et al., 2021). We generated 100 instances for each dataset for the test phase. Appendix E.1 details the data generation process.
Table 1: Evaluation results for combination of $D_{n\times m}$ and $\beta$ with three possible actions (+1, 0, -1). Bold text indicates the best Gap with 100% FR for each combination, while * marks the dataset-wise best among them. “N/A” denotes that calculation is not applicable. FFST for Gurobi is represented as “$<t$ ” because sub-second timing was unavailable in logs.
Method $D_{9\times 18}$ $D_{50\times 20}$ $D_{100\times 50}$ $D_{300\times 100}$
FR $\uparrow$ FFST $\downarrow$ Gap $\downarrow$ FR $\uparrow$ FFST $\downarrow$ Gap $\downarrow$ FR $\uparrow$ FFST $\downarrow$ Gap $\downarrow$ FR $\uparrow$ FFST $\downarrow$ Gap $\downarrow$
Gurobi 100% $<$ 1s 0.00% 100% $<$ 1s 0.00% 100% $<$ 1s 0.01% 100% $<$ 1s 0.03%
$\beta=1$ LS - - 0.07% - - 0.09% - - 0.57% - - 0.78%
M_MLP 100% 0.05s 0.01% 87% 0.06s 0.12% 99% 0.06s 0.52% 97% 0.10s 0.64%
M_CNN 100% 0.06s 0.01% 1% 0.06s 0.05% 0% N/A N/A 0% N/A N/A
Ours 100% 0.05s 0.15% 100% 0.05s 0.09% 100% 0.07s 0.55% 100% 0.13s 0.45%
$\beta=0.6$ LS - - 0.07% - - 0.08% - - 0.58% - - 0.69%
M_MLP 100% 0.05s 0.02% 100% 0.05s 0.12% 99% 0.06s 0.51% 97% 0.10s 0.65%
M_CNN 100% 0.06s 0.00%* 3% 0.04s 0.01% 0% N/A N/A 0% N/A N/A
Ours 100% 0.05s 0.00%* 100% 0.05s 0.08% 100% 0.09s 0.38% 100% 0.11s 0.41%*
$\beta=0.3$ LS - - 0.07% - - 0.08% - - 0.57% - - 0.72%
M_MLP 100% 0.05s 0.00%* 100% 0.05s 0.11% 100% 0.06s 0.40% 99% 0.11s 0.65%
M_CNN 100% 0.06s 0.01% 4% 0.04s 0.03% 0% N/A N/A 0% N/A N/A
Ours 100% 0.05s 0.00%* 100% 0.05s 0.07%* 100% 0.06s 0.34%* 100% 0.11s 0.47%
4.1.2 Compared Methods
For RQ1, we evaluated the RL-MILP solver against the SOTA commercial optimizer, Gurobi, and the classic local search (LS). LS starts in $phase2$ from a feasible solution provided by the trained RL agent in $phase1$, as it requires a feasible starting point (Bertsimas & Tsitsiklis, 1997). For RQ3, we evaluated our method against two baselines: M_MLP and M_CNN. These baselines adopt Multi-layer perceptron (MLP) and Convolutional neural network (CNN), as shown effective in finding a feasible solution (Qi et al., 2021). The MLP and CNN substitute for the Transformer encoder in Figure 5, and they are referred to as M_MLP and M_CNN, respectively. For RQ2, the compared methods were trained with different $\beta$ settings. Existing partial end-to-end learning methods lack support for integer variables and are not standalone solvers; hence, they were excluded. Each method solves a given MILP instance within 100 seconds.
4.1.3 Metrics
For feasibility evaluation, we introduce Feasible rate (FR) and First feasible solution time (FFST). FR measures the ratio of instances where a feasible solution is found. FFST records the time taken to obtain the first feasible solution. For solution quality, we use Relative primal gap (Gap) (Gasse et al., 2022) and #win. Gap quantifies how close a method’s best objective value $obj_{b}$ is to the best-known value. Gap is calculated as $\text{Gap}=\frac{|obj_{b}-BKS|}{|BKS|}*100$, where $BKS$ is the best value obtained among all compared methods, including Gurobi with a 600-second time limit. #win measures the number of test instances where a method reaches the best value obtained among all compared methods. In case of a tie, each method receives a count for #win.
4.2 Main Results
The results in Table 1 show that the RL-MILP solver accurately finds feasible solutions and achieves (near-)optimal solutions. It consistently finds the first feasible solutions within 1 second across all instances and combinations of $D_{n\times m}$ and $\beta$. Additionally, it reaches optimal objective values on $D_{9\times 18}$ and near-optimal values on larger datasets, maintaining a Gap below 1%.
With a toward-optimal bias (i.e., $\beta=\{0.3,0.6\}$), the compared methods tend to find better solutions. The agent chooses the lesser of two evils, as the potential penalty for solutions with $obj_{t+1}<obj_{b}$ is lower than for those with $obj_{t+1}\geq obj_{b}$ (see Appendix B.3). Interestingly, the RL-MILP solver with $\beta=1$ on $D_{9\times 18}$ exhibits the worst Gap among the compared methods, as $\beta=1$ implies no toward-optimal bias. Thus, a suitable toward-optimal bias promotes exploration toward $obj_{r}$, enhancing solution quality. This empirical evidence highlights the importance of guiding the agent’s search, supporting Proposition 1.
Comparing the RL-MILP solver with LS directly highlights search performances in $phase2$. Since LS begins $phase2$ from a feasible solution provided by the RL-MILP solver, feasibility evaluation for LS is omitted. Even on the smallest dataset $D_{9\times 18}$, LS fails to reach the optimal solution, unlike the RL-MILP solver whose neighborhood set is larger.
The RL-MILP solver achieves the lowest Gap with a 100% FR across all datasets, outperforming the baselines, while M_MLP and M_CNN struggle as dataset size increases. These results suggest that the attention mechanism effectively captures variable-constraint interactions and relationships between distant nodes even in larger datasets.
Table 2: Performance by neighborhood sizes on $D_{100\times 50}$. All methods achieves FFST under 1 second.
Method $\Delta_{x}$ $\beta=0.5$ $\beta=0.8$
FR $\uparrow$ Gap $\downarrow$ #win $\uparrow$ FR $\uparrow$ Gap $\downarrow$ #win $\uparrow$
LS 1 100% 0.58% 2 100% 0.58% 0
Ours 1 100% 0.36% 21 100% 0.36% 14
3 100% 0.31% 40 99% 0.32% 37
5 100% 0.31% 50 100% 0.30% 55
4.3 Impact of Varying Neighborhood Sizes
Inspired by the insight that moving farther can overcome barriers created by numerous constraints in the search space (Shaw, 1998), we examine the effect of the agent’s movement magnitude per step on the performance. In this experiment, we extend the solving time to 300 seconds, providing sufficient time for the agent to escape local optima. The neighborhood size grows as the number of changeable variables $n_{x}$ per step and the maximum magnitude of change for variables $\Delta_{x}$ increase. Table 2 shows the impact of varying the neighborhood size. The RL-MILP solver with $\Delta_{x}=1$ achieves a lower Gap than LS with its larger $n_{x}=n$ than LS’s $n_{x}=1$. Furthermore, the highest #win is obtained with $\Delta_{x}=5$, which is the largest action set (+5, +4, +3, +2, +1, 0, -1, -2, -3, -4, -5) in this experiment. This suggests that expanding the neighborhood size enhances the RL-MILP solver’s ability to overcome local optima. As noted in previous research (Son et al., 2023), ensuring the optimality of neural networks for CO problems remains an open challenge due to the lack of theoretical guarantees. This result suggests a promising direction for exploring the optimality of RL-based solvers for MILP problems.
Table 3: Ablation study of the proposed GNN architecture on $D_{100\times 50}$ with $\beta=0.6$ and three possible actions.
Ablation Settings Metric
TE MP PE SL PT FR $\uparrow$ Gap $\downarrow$
100% 0.59%
99% 0.49%
98% 2.82%
99% 2.69%
100% 0.42%
100% 0.38%
4.4 Ablation Studies
We conducted an ablation study to evaluate the effectiveness of the following components: Transformer encoder (TE), MPNN (MP), Periodic embedding (PE), Separated layers (SL), and [PHA] token (PT), illustrated in Figure 5. As shown in Table 3, comparing the removal of TE with the removal of MP, the latter leads improvement in Gap but a loss in FR. Although TE improves Gap, this result emphasizes the importance of MP in capturing local information such as feasibility. Combining all components achieves the best performance in terms of Gap and FR, with SL contributing substantial improvement.
5 Related Work
Bengio et al. (2021) have categorized ML techniques for solving CO into three groups. The first group, Learning to Configure Algorithms, leverages ML to optimize the configuration of specific techniques in traditional solvers. For instance, ML improves solving speed by deciding whether to execute specific operations, such as decomposition (Kruber et al., 2017). The second group, Machine Learning Alongside Optimization Algorithms, integrates ML into traditional solvers to aid key decisions during the optimization process (e.g., selecting the next search region (Lodi & Zarpellon, 2017)). The third group, End-to-End Learning, uses ML to directly learn and predict solutions. This group includes ML-based primal heuristics (Ding et al., 2020; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024), as well as our method. Existing ML-based primal heuristics output only partial solutions and face challenges in handling non-binary integer variables. They also rely on costly supervised learning for labeled data. A previous RL-based study on MILP focuses solely on finding the first feasible solution (Qi et al., 2021) without optimizing further. To our knowledge, this study is the first full end-to-end RL-based solver that improves solution quality for MILP involving integer variables without relying on traditional solvers.
6 Conclusion
In this study, we introduced a novel RL-MILP solver that generates complete feasible solutions without delegating sub-problems to traditional solvers. We designed an RL system that enables the agent to learn MILP problem-solving by capturing variable-constraint relationships. Building on theoretical insights, we guide the RL agent’s solution search by designing a structured reward mechanism and a search strategy. To enhance problem-solving capabilities, we proposed a Transformer-based GNN architecture for the RL-MILP solver. Experimental results show that the RL-MILP solver quickly finds the first feasible solution and improves solution quality toward optimality. The RL-MILP solver achieves optimal solutions for a small dataset and near-optimal solutions with a Gap of less than 1% for larger datasets. This study lays the groundwork for full end-to-end RL-driven approaches to MILP problem-solving.
Acknowledgements
This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. 2019-0-01267, GPU-based Ultrafast Multi-type Graph Database Engine SW).
Appendix A Proof of Proposition 1
In this section, we prove Proposition 1: If a better feasible solution $\mathbf{x}^{\prime}$ exists for a given MILP problem, $\mathbf{x}^{\prime}$ must yield $obj^{\prime}$ that lies between $obj_{r}$ and $obj_{b}$.
A.1 Term Definitions
  • $\mathcal{F}_{MILP}$: The set of feasible solutions (feasible region) for the given original MILP problem.
  • $\mathcal{F}_{LP}$: The set of feasible solutions (feasible region) for the LP-relaxed problem of the original MILP problem.
  • $\mathbf{x}_{r}$: The optimal solution of the LP-relaxed problem.
  • $\mathbf{x}_{opt}$: The optimal solution of the original MILP problem.
  • $\mathbf{x}_{b}$: The current best feasible solution of the original MILP problem.
  • $\mathbf{x}^{\prime}$: A better feasible solution than $\mathbf{x}_{b}$.
  • $obj_{r}$: The optimal value obtained by $\mathbf{x}_{r}$.
  • $obj_{opt}$: The optimal value obtained by $\mathbf{x}_{opt}$.
  • $obj_{b}$: The objective value obtained by $\mathbf{x}_{b}$.
  • $obj^{\prime}$: The objective value obtained by $\mathbf{x}^{\prime}$.
A.2 Background
Adding a constraint to a LP problem introduces a new boundary to the feasible region. The new boundary either reduces the size of the feasible region or leaves it unchanged. As previously mentioned in Section 1, A LP problem becomes a MILP problem if integrality requirements are added. Consequently, a MILP problem has more constraints than its LP-relaxed problem, establishing the relationship $\mathcal{F}_{MILP}\subseteq\mathcal{F}_{LP}$ (Guieu & Chinneck, 1999). Due to this relationship, feasible solutions to the original MILP problem are also feasible for its LP-relaxed problem. Hence, the optimal solution $\mathbf{x}_{opt}$ for the original MILP problem lies within the feasible region of the LP-relaxed problem (i.e., $\mathbf{x}_{opt}\in\mathcal{F}_{LP}$). In the feasible region of the LP-relaxed problem excluding $\mathcal{F}_{MILP}$ (i.e., $\mathcal{F}_{LP}-\mathcal{F}_{MILP}$), there may exist solutions that yield better (i.e., smaller for minimization problems) objective values than $\mathbf{x}_{opt}$. Therefore, in minimization problems, the inequality $obj_{r}\leq obj_{opt}$ holds (Martin, 2012).
A.3 Proof
Assumption 1.
$\mathbf{x}_{b}\neq\mathbf{x}_{opt}$
Under Assumption 1, at least one better feasible solution $\mathbf{x}^{\prime}$ exists. According to the definitions provided in Section A.1, $\mathbf{x}^{\prime}$ represents a better solution than $\mathbf{x}_{b}$. Since we are dealing with a minimization problem, it follows that $obj^{\prime}<obj_{b}$. Additionally, by the definitions in Section A.1, $\mathbf{x}_{opt}$ is the optimal solution for the original MILP problem, which implies $obj_{opt}\leq obj^{\prime}<obj_{b}$. As explained in Section A.2, $obj_{r}\leq obj_{opt}$ holds, establishing that $obj_{r}$ serves as the lower bound for the original MILP problem in a minimization context. Consequently, the relationship $obj_{r}\leq obj_{opt}\leq obj^{\prime}<obj_{b}$ is valid. Hence, since $obj^{\prime}$ derived from $\mathbf{x}^{\prime}$ always falls within the range of $obj_{r}$ and $obj_{b}$, Proposition 1 holds true.
Appendix B Visual Explanation of Reward Function
B.1 Feasibility Reward
The feasibility reward $\mathcal{R}_{t,\text{F}}$ is calculated in proportion to the degree of feasibility improvement or deterioration. In Figure 6, the agent violates Constraint-1 while satisfying Constraint-2. If the agent moves closer to the feasible region, it is considered an improvement in feasibility. Conversely, if it moves further away and ends up violating Constraint-2 as well, it is regarded as a deterioration. In this way, rewards are assigned based on how much the agent’s actions improve or worsen feasibility. As a result, the agent learns to find feasible solutions to maximize its rewards.
Figure 6: Illustration of feasibility improvements and deteriorations.
B.2 Optimization Reward in $phase2$
This section provides a detailed explanation of the reward function in $phase2$ (Eq. 13) with visual aids. If a better objective value than $obj_{b}$ exists, it must lie between the lower bound $obj_{r}$ and $obj_{b}$ (see Appendix A). Thus, the agent should explore solutions within this interval (i.e., $obj_{t+1}\in[obj_{r},obj_{b})$) to obtain a better solution. As $obj_{t+1}$ approaches $obj_{r}$, the objective value improves; however, the feasibility becomes more challenging due to tighter constraints. Therefore, we encourage the agent to explore solutions near the current best feasible solution $obj_{b}$.
[…]
The green circles in Figure 7 correspond to better feasible solutions whose objective values is $obj_{t+1}<obj_{b}$ (Case 1 in Eq. 13). We regard the agent’s actions that fail to find better feasible solutions than $\textbf{x}_{b}$ as incorrect actions. However, feasible solutions outside the interval are neither rewarded nor penalized (Case 2). The reason is not only that such solutions fail to improve solution quality but also that applying penalties could undermine the learning achieved in $phase1$. To maintain consistency with the learning from $phase1$, at least no penalties are imposed on feasible solutions outside the interval. For infeasible solutions, the agent is penalized in proportion to the gap between $obj_{t+1}$ and $obj_{b}$ (Case 3, 4), which corresponds to all triangles. To guide the agent toward the optimal objective value, infeasible solutions with $obj_{t+1}<obj_{b}$ (green triangle) are penalized less than the ones with $obj_{t+1}\geq obj_{b}$ (red triangle) by applying $\beta\in(0,1]$ (Case 3).
Let the objective values be $obj_{r}=-20$, $obj_{b}=-10$, $\mathbf{q_{1}}=-16$, $\mathbf{q_{2}}=-13$, $\mathbf{q_{3}}=-11$, $\mathbf{q_{4}}=-8$, $\mathbf{q_{5}}=-7$, $\alpha=2$, and $\beta=0.5$. For the green circles, rewards in $phase2$ are calculated by the first case in Eq. 13. The reward for $\mathbf{q_{1}}$ is $(1-\frac{|-16-(-20)|}{|-20|})\cdot\alpha=0.8\cdot 2=1.6$. The reward for $\mathbf{q_{3}}$ is $(1-\frac{|-11-(-20)|}{|-20|})\cdot\alpha=0.55\cdot 2=1.1$. Since $\mathbf{q_{1}}$ is closer to the lower bound compared to $\mathbf{q_{3}}$, it receives a higher reward. For the red circle, the reward is 0 by the second case. For the green triangle, the penalty is calculated using the third case. The penalty for $\mathbf{q_{2}}$ is $\mathcal{R}_{t,\text{F}}\cdot\frac{|-13-(-10)|}{|-20-(-10)|}\cdot\beta=\mathcal{R}_{t,\text{F}}\cdot 0.15$. For the red triangle, the penalty is calculated using the fourth case. The penalty for $\mathbf{q_{5}}$ is $\mathcal{R}_{t,\text{F}}\cdot\frac{|-7-(-10)|}{|-20-(-10)|}=\mathcal{R}_{t,\text{F}}\cdot 0.3$. The distances from $obj_{b}$ to $\mathbf{q_{2}}$ and $\mathbf{q_{5}}$ are the same (i.e., $|-13-(-10)|=|-7-(-10)|=3$), but the penalty for $\mathbf{q_{2}}$ is smaller than that for $\mathbf{q_{5}}$ due to scaling by $\beta$.
B.3 Toward-Optimal Bias
Figure 8 visualizes the agent’s potential rewards in $phase2$. The second quadrant represents rewards for solutions with $obj_{t+1}<obj_{b}$, which increase as they approach the lower bound $obj_{r}$. The third and fourth quadrants illustrate penalties that increase with the distance of infeasible solutions from the current best objective value $obj_{b}$, assuming $\mathcal{R}_{t,\text{F}}$ as a constant for simplicity. When $obj_{t+1}$ is the same as $obj_{b}$, the penalty becomes zero, but the agent incurs a penalty of -10 only if the agent takes no action (i.e., $\mathbf{x}_{t}=\mathbf{x}_{t+1}$ in Eq. 15). As illustrated in Figures 8(a) and 8(b), a lower $\beta$ (higher $\alpha$) results in higher potential rewards and reduced penalties for solutions with $obj_{t+1}<obj_{b}$. We can promote the agent to take actions that lead to solutions with $obj_{t+1}<obj_{b}$ by controlling the toward-optimal bias parameters $\alpha$ and $\beta$.
(a) $\beta=1$
(b) $\beta=0.5$
Figure 8: Illustration of the potential rewards of $\mathcal{R}_{t,\text{p2}}$ as a function of the objective value $obj_{t+1}$ in $phase2$. The $x$-axis represents $obj_{t+1}$, and the $y$-axis represents $\mathcal{R}_{t,\text{p2}}$, with the origin set at $obj_{b}$. Assume that $\mathcal{R}_{t,\text{F}}$ remains constant.
Appendix C Pseudo-Code for Learning Algorithm
Algorithm 1 Learning a Policy for RL-MILP Solver using Actor-Critic Algorithm
Input: Agent parameters $\theta$
Parameter: Update limit $N$, Total step limit $T_{max}$, Warm-up step limit $T_{\text{warm}}$, Grace period $T_{\text{grace}}$, Data parameters $\mathcal{P}$
Output: Updated parameters $\theta$
1:  for $N$ updates do
2:     $M\leftarrow\operatorname{GetNewInstance}(\mathcal{P})$See Appendix E.1}
3:     $\mathbf{x}_{0},obj_{r}\leftarrow\operatorname{Relaxation}(M)$
4:     $\mathcal{S}_{0}\leftarrow\operatorname{Observe}(M,\mathbf{x}_{0})$
5:     $\mathbf{x}_{b}\leftarrow\varnothing$;  $obj_{b}\leftarrow\infty$
6:     phase $\leftarrow 1$;  warmup $\leftarrow$ True;  n_steps_grace $\leftarrow 0$
7:     for $t=1,2,\dots,T_{\max}$ do
8:        $obj_{b}^{\prime}\leftarrow obj_{b}$
9:        $\mathcal{A}_{t}\leftarrow\pi_{\theta}(\mathcal{S}_{t})$
10:        $\mathcal{S}_{t+1},\textbf{x}_{b},obj_{b}\leftarrow\operatorname{Search}(\mathcal{S}_{t},\mathcal{A}_{t},\textbf{x}_{b},obj_{b}^{\prime},\texttt{phase})$See Algorithm 2}
11:        $\mathcal{R}_{t,\text{total}}\leftarrow\operatorname{Reward}(M,\mathcal{S}_{t+1},\mathcal{S}_{t},obj_{r},obj_{b}^{\prime},\texttt{phase})$See Equation 6}
12:        $\delta_{td}\leftarrow\mathcal{R}_{t,\text{total}}+\gamma\cdot V_{\theta}(\mathcal{S}_{t+1})-V_{\theta}(\mathcal{S}_{t})$
13:        $\mathcal{L}_{\theta}\leftarrow-\log\pi_{\theta}(\mathcal{A}_{t}\mid\mathcal{S}_{t})\cdot\delta_{td}+\delta_{td}^{2}$
14:        $\theta\leftarrow\operatorname{Update}(\mathcal{L}_{\theta},\theta)$
[…]
17:        else if $\texttt{phase}=2$ then
18:           if $\mathbf{x}_{t+1}$ $\in\mathcal{F}$ then
[…]
20:           else
[…]
22:           end if
23:        end if
[…]
26:           $\textbf{x}_{t+1}\leftarrow\textbf{x}_{b}$
[…]
28:           phase $\leftarrow 2$
[…]
30:        end if
[…]
32:           break
33:        end if
34:     end for
35:  end for
36:  return $\theta$
Appendix D Pseudo-Code for Solution Search Strategy
Our algorithm is designed to search for feasible solutions of a given instance $M$ by exploring the neighborhood of the best solution found so far. The solution search process proceeds as follows: Based on $\mathcal{A}_{t}$, the next observation $\mathcal{S}_{t+1}=(\mathbf{x}_{t+1},\mathbf{f}_{t+1},obj_{t+1})$ are obtained (Line 1). If $\textbf{x}_{t+1}$ is feasible and $obj_{t+1}$ is better (i.e., lower) than or equal to $obj_{b}^{\prime}$, then $obj_{b}$ and $\textbf{x}_{b}$ are updated to the new $obj_{t+1}$ and $\textbf{x}_{t+1}$, respectively (Lines 2-4). Updating $\textbf{x}_{b}$ implies that the center of the local region is also updated. The agent moves back to the center of the local region (i.e., $\textbf{x}_{b}$) if it fails to find a better feasible solution and falls outside the local region in $phase1$ (Lines 5-7). In $phase2$, the agent does not move unless it find a better feasible solution and moves back to the origin (Lines 8-10). The algorithm returns $\mathcal{S}_{t+1}=(\mathbf{x}_{t+1},\mathbf{f}_{t+1},obj_{t+1})$, $\textbf{x}_{b}$, and $obj_{b}$ (Line 12).
Algorithm 2 Solution Search Algorithm
Input: Observation $\mathcal{S}_{t}=(\mathbf{x}_{t},\mathbf{f}_{t},obj_{t})$, Action $\mathcal{A}_{t}$, Best feasible solution $\textbf{x}_{b}$, Best objective value $obj_{b}^{\prime}$,
Current phase phase
Parameter: Local region size $\Delta$
Output: New observation $\mathcal{S}_{t+1}$, Best feasible solution $\textbf{x}_{b}$, Best objective value $obj_{b}$
1:  $\mathcal{S}_{t+1}\leftarrow\operatorname{Move}(\mathcal{S}_{t},\mathcal{A}_{t})$
2:  if $\textbf{x}_{t+1}\in\mathcal{F}$ and $obj_{t+1}\leq obj_{b}^{\prime}$ then
3:     $obj_{b}\leftarrow obj_{t+1}$
4:     $\textbf{x}_{b}\leftarrow\textbf{x}_{t+1}$
5:  else if phase = 1 and $|x_{b,i}-x_{t+1,i}|>\Delta,\,\exists i$ then
6:     $obj_{b}\leftarrow obj_{b}^{\prime}$
7:     $\textbf{x}_{t+1}\leftarrow\textbf{x}_{b}$
8:  else if phase = 2 then
9:     $obj_{b}\leftarrow obj_{b}^{\prime}$
10:     $\textbf{x}_{t+1}\leftarrow\textbf{x}_{b}$
11:  end if
12:  return $\mathcal{S}_{t+1}$, $\textbf{x}_{b}$, $obj_{b}$
Appendix E Detail of Experimental Setup
E.1 Datasets Generation Process
Table 4 summarizes the parameters used for instance generation. Considering that the ratio of non-zero coefficients $\rho$ in typical LP problems is less than 5% (Hillier & Lieberman, 2015), we set a higher density of 10% to promote more interactions between variables in constraints. According to the default settings of the SOTA optimizer Gurobi, the lower bound $l_{i}$ and upper bound $u_{i}$ for decision variables are set to 0 and $\infty$, respectively. Training data is generated on-the-fly whenever the agent completes training on each instance, while test data is pre-generated in advance.
Table 4: Parameters for MILP instance generation.
Parameter Distribution
c $\text{randint}[-10,1]$
A $\text{randint}[1,10]$ with density $\rho=0.1$
b $\textbf{A}\xi+\epsilon$, where
$\xi_{i}\sim\text{randint}[1,10],\forall i=1,\ldots,n$ and
$\epsilon_{j}\sim\text{randint}[1,10],\forall j=1,\ldots,m$
$I$$\{x\mid x\in\mathbb{N},1\leq x\leq n\}$
$l_{i}$ $0,\forall i=1,\ldots,n$
$u_{i}$ $\infty,\forall i=1,\ldots,n$
E.2 Evaluation Environment
We conducted all evaluations under identical configurations. The evaluation machine is equipped with two Intel Xeon Silver 4214R CPUs @ 2.4GHz, 768GB RAM, and eight NVIDIA Titan RTX GPUs. All experiments were performed using a single NVIDIA Titan RTX GPU. The software environment includes PyTorch 1.10.2, Gymnasium 0.29.1, and Gurobi 9.5.2.
E.3 Method Implementation
We utilize the bipartite graph convolution available on GitHub [^1] https://github.com/sribdcn/Predict-and-Search_MILP_method (Han et al., 2023), as the architecture for our MPNN. Two iterations of the process shown in Figure 2(a) are applied, resulting in two constraint-side and two variable-side convolutions. Our proposed model is implemented using the Transformer encoder code from GitHub [^2] https://github.com/ucbrise/graphtrans (Wu et al., 2021), maintaining the same configuration. We developed two MPNN-based baselines, M_MLP and M_CNN. M_MLP consists of four MLP layers with a hidden size of 128 and tanh activation, while M_CNN includes four CNN layers followed by an MLP layer with ReLU activation. We utilized the positional encoding module from GitHub [^3] https://github.com/yandex-research/rtdl-num-embeddings (Gorishniy et al., 2022).
All ML models were trained using the proposed learning algorithm (Algorithm 1) with RMSprop (learning rate = 1e-4, epsilon = 1e-5, alpha = 0.99, weight decay = 1e-3). They were trained concurrently on 64 different instances with 5,000 parameter updates for the results in Tables 1 and 3, and 10,000 for Table 2. Our RL algorithm is built upon the Actor-Critic implementation in PyTorch [^4] https://github.com/ikostrikov/pytorch-a2c-ppo-acktr-gail (Kostrikov, 2018), modified to be tailored for MILP problems.